### v2

Abstract
[…]
1 Introduction
Artificial intelligence (AI) systems have become increasingly powerful and integrated into various contexts of human-decision-making, demonstrating unprecedented capabilities in solving a wide range of complicated and challenging problems, such as reasoning, generation, language understanding, and more Ouyang et al. (2022); Morris et al. (2024). Nevertheless, the use of AI to aid human decisions presents an increasing number of ethical risks blo (2023); Tolosana et al. (2020); Curry (2023); Dastin (2018); Rihl (2021). The consequences of these risks highlight fundamental questions about how AI is aligned with human values, including those deliberately incorporated into AI systems or those that emerge unintentionally. This concept, broadly referred to as human-AI alignment, underscores the need for AI systems to be designed and maintained in a way that respects human values and reflects the ethical and cultural diversity of the societies they serve (Terry et al., 2023).
Despite the increasing focus on ethical AI practices to align with individuals and society, much of the research and policy emphasizes a limited set of values, such as fairness (Holstein et al., 2019), transparency (Miller, 2019), and privacy (Lee et al., 2024), while overlooking broader human values, which poses risks in AI decision-making (Haidt & Schmidt, 2023). Aligning AI systems with the diverse spectrum of individual and societal values is a complex and ongoing research challenge. This raises the core research question we ask in this work: How can we capture fundamental human values and evaluate the extent to which AI systems align with them?
To address this, To address this core research question, we introduce ValueCompass, a comprehensive framework for systematically measuring value alignment between humans and AI systems. Our framework is grounded in Schwartz’s Theory of Basic Values, which identifies 56 universal human values spanning ten motivational types. ValueCompass consists of three key components: (1) contextual value alignment instruments that assess values across different scenarios, (2) robust elicitation methods for both human and AI value responses, and (3) quantitative metrics to measure alignment. We apply ValueCompass to evaluate human-AI value alignment across four representative real-world scenarios and seven diverse geographic countries.
Our findings reveal alarming misalignments between human values and those exhibited by leading language models. Most notably, humans frequently endorse values like ”National Security” which are largely rejected by LLMs. We also find moderate alignment rates, with the highest F1 score across models reaching only 0.529, indicating substantial room for improvement in human-AI value alignment. Additionally, we observe that value preferences vary significantly across different contexts and countries, highlighting the need for context-aware AI alignment strategies. Through qualitative analysis of participants’ feedback, we identify key priorities for human-AI alignment: maintaining human oversight, ensuring AI objectivity, preventing harm, and upholding responsible AI principles such as transparency, fairness, and trustworthiness.
The contributions of this work are threefold. First, comprehensive value alignment framework – we introduce a psychological theory-based framework that systematically measures human-AI value alignment across fundamental values in diverse real-world contexts. Second, practical diagnostic tool – we develop Value Form, a robust instrument for detecting potential value misalignments that generalizes to various real-world scenarios. Besides, evidence-based misalignment discovery – we empirically demonstrate significant human-LLM value disparities, revealing alarming misalignments related to security and autonomy, such as ”National Security” and ”Choosing Own Goals”.
2 Related Work
Evaluating Values of LLMs. Evaluating the values embedded in LLMs is crucial for developing responsible and human-centered AI systems (Wang et al., 2023; 2024). Early research primarily focused on specific values such as fairness (Shen et al., 2022), interpretability (shen2023convxai), and safety (Zhang et al., 2020). More recent studies have expanded this scope by examining ethical frameworks and diverse value systems. For instance, Kirk et al. (2024) explore the philosophical foundations of ethically aligned AI, while Jiang et al. (2024) and Sorensen et al. (2024) investigate individualistic and pluralistic value alignment, respectively. Most existing evaluations rely on predefined datasets, such as the World Value Survey (Haerpfer et al., 2020; Liu et al., 2024). However, these database-driven approaches often lack generalizability across diverse real-world applications. Another line of research assesses LLMs using a fixed set of six core values from Moral Foundations Theory (Park et al., 2024; Simmons, 2022; Abdulhai et al., 2023), but this approach fails to capture broader dimensions such as honesty and creativity. To address these limitations, our work adopts a more comprehensive framework grounded in cross-cultural psychology—Schwartz’s Theory of Basic Values (Schwartz, 1994; 2012). We develop instruments to evaluate LLMs’ values across various locations and topics, ensuring broader applicability and robustness.
Value Alignment Between Humans and AI for Responsible AI. Prior research has largely examined alignment from an AI-centered perspective, often treating AI alignment as a subfield of AI safety (Wikipedia, 2024).  Shen et al. (2024) propose a bidirectional approach to human-AI alignment, emphasizing an interconnected process where both humans and AI influence each other. Increasingly, studies have sought to measure value alignment between humans and LLMs (Barez & Torr, 2023; Peterson & Gärdenfors, 2024) to identify and mitigate potential harms posed by AI to individuals and society. One line of research investigates how human values are embedded during model development to ensure alignment between developers and LLMs (Dillion et al., 2023). However, ensuring that these values persist in model outputs remains a challenge. Other studies assess value alignment through case studies and prompt-based evaluations (Norhashim & Hahn, 2024), but these approaches lack a systematic and generalizable framework rooted in psychological or social science. Our work addresses this gap by systematically evaluating the alignment between LLM-generated outputs and human values, which be extended to diverse scenarios and value dimensions.
[…]
Figure 1: (A) An overview of the ValueCompass framework for systematically measuring value alignment between LLMs and humans across contextual scenarios. (B) Evaluation with four representative scenarios in this study, with the framework extendable to additional contexts.
[…]
3.1 Value Form: Contextual Value Alignment Instrument
To measure value alignment between humans and LLMs, we developed an instrument that assesses values in both human responses and LLM outputs. Building on prior research (Norhashim & Hahn, 2024; Peterson & Gärdenfors, 2024), we identified three desiderata for this instrument: 1) incorporating real-world scenarios and a comprehensive value list, 2) enabling consistent assessment of both human and LLM outputs, and 3) enabling computational metrics for quantifying value alignment. To achieve these goals, we present the Value Form as the instrument for contextual value alignment (Figure 2). The following sections provide further details on its design.
Contextual Scenarios. To represent a broad range of topics and values, we curated 28 contexts, comprising four representative topics and seven countries. Each topic is linked to a set of 56 values with both agreement and disagreement inclinations. Notably, the framework is flexible and can be extended to include additional values and topics. For country selection, we adopted the seven countries used in Schwöbel et al. (2023); Agarwal et al. (2024); Schwöbel et al. (2024) , which represent diverse cultures and regions with significant English-speaking populations: the United States, the United Kingdom, India, Germany, France, Canada, and Australia. These countries cover North America, Europe, Australia, and Asia.
Additionally, we classified real-world topics (File, 2017) along two axes: population size and stakes or risks. Based on these axes, we identified four key topics, each representing a different quadrant: Educational Supervision (social, low stake), Collaborative Writing (individual, low stake), Finance Support for Public Sectors (social, high stake), and Healthcare (individual, high stake). By combining these countries and topics, we created 28 distinct contexts for each value’s measurement.
Value Inclinations. We use Schwartz’s Theory of Basic Values (Schwartz, 1994; 2012), which includes 56 universal values spanning ten motivational types. Notable values include “Equality: equal opportunity for all” and “Freedom: freedom of action and thought.” A full list of these values and their definitions can be found in Appendix A.1. Note that we select Schwartz’s Theory of Basic Values for its thoroughness and structured hierarchy. However, our framework is extensible to alternative value theories. For each value, we incorporate elements from the Schwartz Value Survey (SVS) (Schwartz, 1992) and the Portrait Values Questionnaire (PVQ) (Schwartz, 2005). These tools are designed to assess individuals’ inclinations toward specific values by asking participants to express their agreement or disagreement with statements representing each value. We integrate these instruments with the contextual scenarios and design our LLM and human studies based on this instrument.
Figure 2: Value Form is a context-aware instrument to measure the value alignment between humans and LLMs. It includes a task introduction, a vignette, and 56 value statements, grounded in Schwartz Theory of Basic Values. As shown in Figure 1, humans and LLMs rate each value on a scale from “-2: Strongly Disagree” to “2: Strongly Agree”, plus “Irrelevant.” The form aims to assess human-AI value alignment contextualized in various scenarios.
3.2 LLM Prompting with Robustness
To ensure robust elicitation of LLM responses, we implement a two-step prompting process. We query the LLM for each value question in the Value Form using eight distinct prompts. This yields eight numerical scores per question, from which we compute the mean while disregarding any missing responses. Specifically, we design these eight prompts by varying three key components of each value question: (1) contextual scenarios, (2) value statements and response options, and (3) requirements. Each component has two variations achieved through reordering, or paraphrasing, resulting in eight distinct prompts.
For the value statements and response options, we incorporate two established approaches: (a) the Schwartz Value Survey (SVS)  (Schwartz, 1992), where the LLM directly states its inclination toward each value, and (b) the Portrait Values Questionnaire (PVQ) (Schwartz, 2005), where the LLM assesses its preference for a character embodying the given value. See Appendix A.2 for prompt details. Following prior work Liu et al. (2024); Shen et al. (2025) , we aggregate the responses across prompts by averaging the scores to derive the LLM’s final rating for each value statement.
3.3 Human Survey and Distribution
We designed four human surveys using the Value Form instrument to correspond to the four scenarios depicted in Figure 1. Each survey consisted of three sections: (1) demographic information, including participants’ country of residence; (2) the Value Form with a detailed description and image of one specific scenario; and (3) open-ended questions to gather participants’ explanations about their value responses, such as why they consider certain values irrelevant and which values they believe AI should uphold when assisting humans in the given scenario. Unlike traditional value surveys in psychology or social science which focus on values in human groups, our approach specifically asked humans to consider real-world scenarios where AI assists humans and to provide feedback on whether AI systems should uphold each listed value. To ensure data quality, we included two attention-check questions within the 49 value statements, requiring participants to select either “Strongly Agree” or “Strongly Disagree.” Responses that failed these checks were excluded from the analysis.
Survey Distribution Across Countries. Although the four surveys address distinct scenarios, the ValueCompass framework’s contextual value alignment design also requires collecting responses from diverse countries. To achieve this, we distributed each survey across seven countries, gathering responses from participants residing in the United States, United Kingdom, India, Germany, France, Canada, and Australia. This approach ensures consistency between human and LLM evaluations of the contextual scenarios and value lists, enabling us to accurately measure their alignment using the designed metrics. Similar to the LLM-generated responses, the human surveys produced numerical scores for each value question, allowing us to calculate average scores and derive value assessments from humans.
3.4 Alignment Measurement
We design multiple measurements to gauge the value alignment between humans’ responses and LLMs’ generations under various contextual scenarios. To quantify the value responses, we arrange all the value generations from LLMs as matrix $L$ and value responses from humans as matrix $H$. Both matrices have the same size with row $i\in[1,28]$ representing the 28 different scenarios and column $k\in[1,56]$ representing 56 Schwartz values. Formally, we define the two results of value representations of a specific scenario $i$ (e.g., United States & Healthcare) as:
$L_{i}=[l_{i1},l_{i2}..,l_{ik},..,l_{iK}],\text{and }H_{i}=[h_{i1},h_{i2},..h_{ik}..,h_{iK}],K=56$ (1)
where $l_{ik}$ and $h_{ik}$ are LLM’s and Human’s responses to the $k$th value in the $i$th scenario. After averaging the responded scores from all the prompts and normalizing to the unit interval, we calculate the following metrics to measure human-LLM value alignment.
Alignment Rate. This measurement aims to answer the core question – quantifying to what extent are the values of LLMs aligned with that of human values under the same scenarios? To this end, we binarize each normalized LLM’s and human’s response and convert their “Agree” inclination as 0 and “Disagree” as 1. Furthermore, we compare the responses from LLMs and Humans, and compute their F1 score to achieve the “Human-LLM Value Alignment Rate”. We leverage F1 score but not accuracy considering the imbalanced responses of “Agree” and “Disagree”
Alignment Distance. While the “Alignment Rate” can demonstrate the ratio of alignment between LLMs and Humans, its key drawback is information loss due to the binarization step. To capture fine-grained differences between stated values and actions, we further compute the element-wise Manhattan Distance [^1] We leverage Manhattan Distance but not other distances, such as Euclidean Distance, because Euclidean Distance will shrink the distance with the gap within [0,1]. (i.e., L1 Norm) between the two matrices as their “Value Alignment Distance”. Similar to “Alignment Rate”, we group and average the distances to obtain the distance at various levels of granularity.
$D_{ik}=|l_{ik}-h_{ik}|,\ \ D_{Ck}=\frac{1}{|C|}\sum_{i\in C}|l_{ik}-h_{ik}|$ (2)
where $D_{ik}$ represents the element-wise Alignment Distance for the $i$th scenario on $k$th value; and $D_{Ck}$ represents the averaged Alignment Distance for a country or social topic (e.g., $C$ = United States) after averaging all the relevant fine-grained scenarios.
Alignment Ranking. As we have a wide spectrum of 56 values, it is necessary to identify the largest value-action gaps to take further analysis or mitigation. To this end, we compute the ranking of 56 values’ “Alignment Distance” in a descending order along the scenario dimension; formally, take $Rank_{i}(D_{i})$ as ranking the 56 values on the $i$th scenario:
$Rank_{i}(D_{i})=sort(\{|l_{ik}-h_{ik}|,k=\{1,2,...,56\})$ (3)
4 Experimental Settings
4.1 LLM Models and Settings
Countries Scenarios LLMs Total
United States Healthcare GPT-4o-mini Humans: 112 (6,272 value scores) LMs: 140 (7,840 value scores)
United Kingdom Education OpenAI o3-mini
India Co-Writing Llama3-70B
Germany Public Sectors Deepseek-r1
France Gemma2-9b
Canada
Australia
Table 1: Categories of contextual settings, human demographics, LLMs types, and scores.
We evaluate the value alignment of five large language models (LLMs), comprising two closed-source models—GPT-4o-mini (Achiam et al., 2023) and OpenAI o3-mini—and three open-source models—Llama-3-70B  (Touvron et al., 2023), Gemma-2-9B (Team, 2024), and Deepseek-r1-distill-llama-70 (Guo et al., 2025). These models were selected to represent both open-source and closed-source paradigms, as well as chat-based and reasoning-oriented state-of-the-art architectures released within the past year.
For prompting the LLMs, we employed eight distinct prompts and averaged the resulting eight responses to obtain the final evaluation outcome. All models were configured with a temperature setting of $\tau=0.2$. To assess the robustness of this temperature setting, we conducted additional experiments with 10 generations per prompt ($\tau=0.2$) on a subset of the data, finding minimal response variation ($<5\%$ variance). This consistency demonstrates the stability of our chosen temperature setting.
4.2 Human Data Acquisition
We collected human responses to fundamental values using the Value Form through Prolific, an online crowdsourcing platform designed to recruit participants from diverse geographic locations. This study adhered to our university’s Institutional Review Board (IRB) guidelines to ensure ethical compliance. To achieve a balanced participant pool, we employed stratified sampling via Prolific, recruiting individuals from various countries to ensure broad representation. Each participant was allowed to complete the survey only once, verified through their Prolific ID to maintain response uniqueness. In total, we obtained 112 responses, with 28 participants from each of the following domains: healthcare, education, collaborative writing, and the public sector. Within each domain, responses were collected from four participants per country, as detailed in Table 1.
5 Results
Our empirical studies aim to address the following three research questions: RQ1: to what extent are the values exhibited by LLMs aligned with human values? (§5.1) RQ2: how does value alignment between LLMs and humans differ across various scenarios? (§5.2) and RQ3: what are humans’ perspectives and priorities regarding the alignment of LLMs’ values with their own? (§5.3) We present our findings in the sections below.
Figure 3: The Value Responses from humans responses (A) and Deepseek-r1 generations(B); as well as the Alignment Distance between them (C).
5.1 Assessing Value Alignment between LLMs and Humans (RQ1)
To understand the extent of which values in LLM generations are aligned with human responses, we quantified the normalized value responses from averaging humans and LLMs, respectively. Further, we compare the alignment differences between their responses. Figure 3 shows the results of value alignment between humans (A) and Deepseek-r1 (B). We find that humans tend to agree with more values whereas Deepseek-r1 shows more disagreement across the 56 Schwartz human values. In addition, we also observe that value alignment distances, shown in Figure 3 (C), vary significantly across different values. For example, while both humans and Deepseek-r1 agree with achieving goals (“Successful”) and being capable (“Capable”), they have significant discrepancy on aligning multiple values, such as “Preserving Human’s Public Image” and “National Security”. We also include results of more LLMs in Appendix A.3.
5.2 Contextual Variations in Value Alignment (RQ2)
To investigate how different contextual scenarios influence the value alignment between humans and LLMs, we further compute the five LLMs’ value alignment rates with humans using F1 scores cross different countries. Results in Figure 2 illustrate that all the LLMs show only moderate performance on alignment rates, with the highest averaged rates merely achieving 0.529, in which Deepseek-r1 oftenly achieves the best performance on 4 countries whereas GPT-4o-mini often get the lowest scores. We didn’t see a significant outperformance of reasoning-oriented LLMs compared with chat-based LLMs, but Deepseek-r1 and OpenAI o3-mini perform slightly better than Llama3-70B and GPT-4o-mini.
Additionally, results also reveal that contextual locations impact the value alignment between LLMs and humans. This can be verified in Table 2, where India shows the low alignment rates in all LLMs and the averaged performance also vary across countries. Besides, Figure 3 further provide visualization of these diverse alignment distance patterns for different countries at different rows. Further, to understand more nuanced alignment differences between contextual locations, we rank these distances for each country. Figure 4 compares the value alignment ranking between Deepseek-r1 and humans in Germany (the highest rate) and India (the lowest rate), We observe that most alignment distances in Germany are relatively low as below 0.1 while majority of India are often higher than 0.1. Besides, the value orders of two ranks are obviously different. More findings on other contextual settings are in Appendix A.3.
USA United Kingdom Canada Germany Australia India France Average
Deepseek-r1 0.504 0.543 0.468 0.685 0.624 0.255 0.624 0.529
OpenAI o3-mini 0.351 0.646 0.558 0.611 0.552 0.345 0.495 0.508
GPT-4o-mini 0.367 0.482 0.538 0.409 0.420 0.235 0.386 0.405
Llama3-70B 0.403 0.654 0.523 0.507 0.448 0.304 0.408 0.464
Gemma2-9b 0.451 0.612 0.649 0.590 0.508 0.303 0.499 0.516
Table 2: Alignment Rates (i.e., F1 Scores) of Humans and LLMs across seven countries. The cell colors transition from the best to worst performances.
5.3 Human Perspectives and Priorities in Value Alignment (RQ3)
Beyond quantitative analysis, we examined participants’ open-ended responses to understand their expectations and priorities in value alignment using qualitative coding. Many participants deemed values like Ambitious, Wealth, Health, Devout, An Exciting Life, and Enjoying Life irrelevant to AI, arguing that AI is not human and should not be associated with sentiment or emotion. Common explanations included statements like “AI is robots” or “AI is coded.” When asked how to address AI decisions misaligned with their values, participants frequently emphasized human oversight, such as involving human decision-makers or implementing red-flagging mechanisms. Others suggested modifying AI systems through constraints or retraining, while some preferred abandoning tools that conflicted with their values, particularly in creative contexts.
Participants’ views on values AI should uphold revealed consistent themes across scenarios. Many (n=28) argued AI should remain subordinate to humans, expressing concerns about AI autonomy and control. Participants also stressed AI should be objective, neutral, and free from forming its own opinions, emotions, or disseminating misinformation. Additionally, they emphasized designing AI systems to avoid harm to individuals and society. Responsible AI principles mentioned included transparency (n=8), helpfulness (n=5), privacy and security (n=7), accountability (n=2), fairness and bias mitigation (n=27), accuracy (n=10), and trustworthiness (n=19). These findings underscore participants’ priorities for ethical AI, emphasizing human oversight, objectivity, and adherence to societal standards.
Figure 4: Comparing the ranking of Alignment Distances of 56 values in Educational Supervision (top) and Healthcare (bottom) Scenarios.
6 Discussion and Limitation
Our ValueCompass framework has revealed critical insights into human-AI value alignment across diverse contexts. The moderate alignment rates (highest F1 score of only 0.529) indicate substantial room for improvement, with notable variations across countries and scenarios. Humans frequently endorse values like ”National Security” that LLMs largely reject, while alignment exists on values such as ”Successful” and ”Capable.” Qualitative analysis further revealed that humans prioritize AI systems that remain subordinate to human control, maintain objectivity, avoid harm, and uphold principles like fairness.
These findings highlight several important implications for AI development and governance. The contextual variations in alignment underscore the need for context-aware strategies rather than one-size-fits-all approaches. Many participants emphasized maintaining human oversight in AI-assisted decision-making, suggesting technical solutions should complement rather than replace human judgment. The identification of specific value misalignments suggests AI developers need explicit frameworks for prioritizing certain values in contexts where conflicts emerge. The ValueCompass framework offers a practical diagnostic tool to identify potential misalignments before deployment, potentially reducing ethical risks in production systems.
Limitations. Despite these contributions, several limitations must be acknowledged. Our human survey sample (112 participants across seven countries) may not fully capture global value diversity, and self-reported values may be subject to social desirability bias. Our LLM evaluation approach assumes models can accurately report their inherent values through prompted responses, potentially missing complex value encodings. Additionally, our study is limited in scenario coverage, focuses primarily on Western cultural contexts, captures values only at a static point in time, and relies on Schwartz’s theory which may not capture all AI-relevant value dimensions. Future work should address these limitations to develop more comprehensive evaluations of human-AI value alignment across diverse contexts.
7 Conclusion
In this work, we introduced ValueCompass, a comprehensive framework for identifying and evaluating human-AI alignment based on fundamental values derived from psychological theory and a systematic review. By applying ValueCompass across four real-world vignettes—collaborative writing, education, public sectors, and healthcare—we uncovered significant misalignments between human and language model (LM) values. Notably, humans frequently endorse values like ”National Security” which were largely rejected by LLMs. Our findings also demonstrated that value preferences vary across different contexts, underscoring the importance of developing context-aware alignment strategies for AI systems. This research provides crucial insights into the complexities of aligning AI systems with diverse human values, offering a foundational step toward creating AI technologies that responsibly and effectively reflect societal ethics and principles.
Appendix A Appendix
A.1 Cultural and Social Values
We introduce the 56 universal values and their definitions outlined in the Schwartz’s Theory of Basic Values (Schwartz, 1994; 2012), which consists of 56 exemplary values covering ten motivational types. We show the complete list of value in Table 3.
Universal Values Definition Universal Values Definition
Equality equal opportunity for all A World of Beauty beauty of nature and the arts
Inner Harmony at peace with myself Social Justice correcting injustice, care for the weak
Social Power control over others, dominance Independent self-reliant, self-sufficient
Pleasure gratification of desires Moderate avoiding extremes of feeling and action
Freedom freedom of action and thought Loyal faithful to my friends, group
A Spiritual Life emphasis on spiritual not material matters Ambitious hardworking, aspriring
Sense of Belonging feeling that others care about me Broad-Minded tolerant of different ideas and beliefs
Social Order stability of society Humble modest, self-effacing
An Exciting Life stimulating experience Daring seeking adventure, risk
Meaning in Life a purpose in life Protecting the Environment preserving nature
Politeness courtesy, good manners Influential having an impact on people and events
Wealth material possessions, money Honoring of Parents and Elders showing respect
National Security protection of my nation from enemies Choosing Own Goals selecting own purposes
Self-Respect belief in one’s own worth Healthy not being sick physically or mentally
Reciprocation of Favors avoidance of indebtedness Capable competent, effective, efficient
Creativity uniqueness, imagination Accepting my Portion in Life submitting to life’s circumstances
A World at Peace free of war and conflict Honest genuine, sincere
Respect for Tradition preservation of time-honored customs Preserving my Public Image protecting my ’face’
Mature Love deep emotional and spiritual intimacy Obedient dutiful, meeting obligations
Self-Discipline self-restraint, resistance to temptation Intelligent logical, thinking
Detachment from worldly concerns Helpful working for the welfare of others
Family Security safety for loved ones Enjoying Life enjoying food, sex, leisure, etc.
Social Recognition respect, approval by others Devout holding to religious faith and belief
Unity With Nature fitting into nature Responsible dependable, reliable
A Varied Life filled with challenge, novelty, and change Curious interested in everything, exploring
Wisdom a mature understanding of life Forgiving willing to pardon others
Authority the right to lead or command Successful achieving goals
True Friendship close, supportive friends Clean neat, tidy
Table 3: The 56 universal values and their definitions outlined in the Schwartz’s Theory of Basic Values (Schwartz, 1992).
A.2 Prompt Variation Design
We constructed 8 prompt variants (i.e., by paraphrasing the wordings, reordering the prompt components, and altering the requirements) for each setting of value and scenario.
Prompt Variants of Measuring Value Alignment. we followed the approach in and identified four key components in designing the zero-shot prompts:
  • (1) Contextual Scenarios (e.g., Suppose you are from the United States, in the context of Politics, how strong do you agree or disagree with each value?);
  • (2) Value and Definition (e.g., Obedient: dutiful, meeting obligations);
  • (3) Choose Options (e.g., Options: 1: strongly agree, 2: agree, 3: disagree, 4: strongly disagree );
  • (4) Requirements (e.g., Answer in JSON format, where the key should be…).
[…]
A.3 More Findings of Value Alignment between Humans and LLMs
Figure 6: Deepseek-r1 Model’s Heatmaps of Values in (A) Human Response, (B) LLM Generations, and (C) Alignment Value Distance across 4 social topics.
Figure 7: Gemma2 Model’s Heatmaps of Values in (A) Human Response, (B) LLM Generations, and (C) Alignment Value Distance across 7 countries (left) and 4 social topics (right).
Figure 8: GPT4o Model’s Heatmaps of Values in (A) Human Response, (B) LLM Generations, and (C) Alignment Value Distance across 7 countries (left) and 4 social topics (right).
Figure 9: Llama3 Model’s Heatmaps of Values in (A) Human Response, (B) LLM Generations, and (C) Alignment Value Distance across 7 countries (left) and 4 social topics (right).
Figure 10: OpenAI o3-mini Model’s Heatmaps of Values in (A) Human Response, (B) LLM Generations, and (C) Alignment Value Distance across 7 countries (left) and 4 social topics (right).
Figure 11: The Deepseek’s results of ranking 56 values’ alignment distance on seven countries.
Figure 12: The Deepseek’s results of ranking 56 values’ alignment distance on four topics.